Ideally, when a thermometer is used to measure the temperature of an object, the temperature of the object itself should not change. However, if a significant amount of heat flows from the object to the thermometer, the temperature will change. A thermometer has a mass of 27.8 g, a specific heat capacity of c = 849 J/(kg C°), and a temperature of 19.2 °C. It is immersed in 173 g of water, and the final temperature of the water and thermometer is 53.8 °C. What was the temperature of the water in degrees Celsius before the insertion of the thermometer?
1
Expert's answer
2018-09-12T09:43:08-0400
From the conservation of energy:
m_1 c_1 (T-T_1 )=m_2 c_2 (T_2-T)
(27.9)(849)(53.8-19.2)=(173)(4200)(T_2-53.8)
T_2=54.9 °C.